# Cascades in transport and optical conductivity of Twisted Bilayer Graphene

### v1

@@ Abstract
Among the plethora of correlated states in magic angle TBG, the cascades in the Density of States (DOS) [1, 2, 3, 4] and in the inverse compressibility [5, 6, 7], and the resistive states [8, 9, 6, 10, 11, 12], are the signatures of the electronic correlations that remain up to the highest temperatures. Experimentally, maxima in the resistivity as a function of doping appear around integer fillings up to temperatures much higher than those at which insulating behavior (resistivity increasing as the temperature decreases) is observed and, in many cases, even when the latter is absent  [8, 9, 6, 10, 11, 12]. Evidence of symmetry breaking order has been reported only up to a few kelvin. In many strongly correlated materials, Mott-like physics produces a redistribution of the spectral weight, even if the correlated material is in its normal state, above any symmetry-breaking ordering, or it is a metal [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. However, due to the intrinsic topology of the TBG flat bands, most explanations of the correlated states observed in these systems, including those resilient with temperature, involved the assumption of symmetry breaking [2, 3, 1, 5, 23, 24, 25, 26, 27].
@@
Optical spectroscopy has played a pivotal role in identifying the signatures of electronic correlations in many materials [14]. Optical transitions conserve spin, valley and momentum, making optical probes an ideal set-up to explore the spectrum of moiré heterostructures [35, 36, 37, 38, 39, 40, LiArXiv2024, 41, 42]. At low frequencies, it provides information on the scattering rate of the carriers and their contribution to transport. Due to experimental challenges, doping dependent optical measurements of TBG are not available yet but they could be soon achieved [39, LiArXiv2024, 41, 42].
@@
We focus on a slightly particle-hole asymmetric TBG with twist angle $\theta=1.08^{\circ}$ and interlayer tunneling ratio between the AA and AB regions of the moiré unit cell $w_{0}/w_{1}=$0.7, Fig. 1(g) [43, 44, 45]. We fit the band structure with an 8 orbital model per valley and spin [46, 28, 19, 43] with topologically fragile flat bands. To calculate the interactions between the orbitals in the effective model, we assume a $1/r$ interaction between the electrons at the carbon atoms and relative dielectric constant $\epsilon=16$ [43]. We distinguish two types of orbitals: local AAp and itinerant $lc$. The two AAp orbitals, with $p_{+}$ and $p_{-}$ symmetries, centered at the AA region of the unit cell account for the spectral weight of the flat bands everywhere except close to the $\Gamma$ point. The intra-unit cell interaction between these local AAp orbitals, $U\sim$ 32.6 meV, is much larger than their bandwidth. These orbitals are strongly correlated [28, 19] and tend to form local moments. The $lc$ orbitals are less correlated. The model resembles an extended heavy-fermion like model with a strongly ${\bf k}$-dependent hybridization between the local AAp and the itinerant $lc$ orbitals, concentrated around $\Gamma$, see Fig. S1 [43] and [18, 28, 29, 19, 47, 48, 21, 49]. The interactions involving the itinerant orbitals are of order $U/2$ and cannot be neglected [43, 28, 19, 1, 29]. As in our previous work, we treat the intra and inter-orbital onsite interaction $U$ between the AAp orbitals within DMFT and the other interactions at a mean field Hartree level, within a double self-consistency loop (DMFT+H), see [19, 43]. Temperature is $5.8\,\rm K$ except otherwise indicated. Spontaneous symmetry-breaking is not allowed in the calculation.
@@
Consistent with our previous work [19] and with experimental results [1, 5], the DOS and the inverse compressibility in Fig. 1(a)-(b), respectively, show cascades of spectral weight and asymmetric saw-tooth peaks resetting at integer values of doping. Regions of negative inverse compressibility appear below the integer values after the contribution of the gate is taken into account [43, 20], in agreement with experimental data. The ratio between the AAp onsite interaction $U$, and $\Delta$, the energy gap between the remote and the flat bands at $\Gamma$, $U/\Delta=$0.85, is smaller than the one used in our previous calculations ($U/\Delta=2$ in [19]) and in other works [34]. This prevents the remote bands from crossing the chemical potential at dopings $|\nu|<4$ due to the strong Hartree deformation of the bands at $\Gamma$ , see Fig. 3(d)-(e) and [43]. Consequently, the system is insulating at $|\nu|=4$, instead of metallic as in [19, 34].

### v2

@@ Abstract
Among the plethora of correlated states in magic angle TBG, the cascades in the Density of States (DOS) [1, 2, 3, 4] and in the inverse compressibility [5, 6, 7], and the resistive states [8, 9, 6, 10, 11, 12], are the signatures of the electronic correlations that remain up to the highest temperatures. Experimentally, maxima in the resistivity as a function of doping appear around integer fillings up to temperatures much higher than those at which insulating behavior (resistivity increasing as the temperature decreases) is observed and, in many cases, even when the latter is absent  [8, 9, 6, 10, 11, 12]. Evidence of symmetry breaking order has been reported only up to a few kelvin. In many strongly correlated materials, Mott-like physics produces a redistribution of the spectral weight, even if the correlated material is in its normal state, above any symmetry-breaking ordering, or it is a metal [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. However, due to the intrinsic topology of the TBG flat bands, most explanations of the correlated states observed in these systems, including those resilient with temperature, involved the assumption of symmetry breaking [2, 1, 5, 23, 24, 25, 26, 27].
@@
Optical spectroscopy has played a pivotal role in identifying the signatures of electronic correlations in many materials [14]. Optical transitions conserve spin, valley and momentum, making optical probes an ideal set-up to explore the spectrum of moiré heterostructures [35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. At low frequencies, it provides information on the scattering rate of the carriers and their contribution to transport. Due to experimental challenges, doping dependent optical measurements of TBG are not available yet but they could be soon achieved [39, 42, 43, 44].
@@
We focus on a slightly particle-hole asymmetric TBG with twist angle $\theta=1.08^{\circ}$ and interlayer tunneling ratio between the AA and AB regions of the moiré unit cell $w_{0}/w_{1}=$0.7, Fig. 1(g) [45, 46, 47]. We fit the band structure with an 8 orbital model per valley and spin [48, 28, 19, 45] with topologically fragile flat bands. To calculate the interactions between the orbitals in the effective model, we assume a $1/r$ interaction between the electrons at the carbon atoms and relative dielectric constant $\epsilon=16$ [45]. We distinguish two types of orbitals: local AAp and itinerant $lc$. The two AAp orbitals, with $p_{+}$ and $p_{-}$ symmetries, centered at the AA region of the unit cell account for the spectral weight of the flat bands everywhere except close to the $\Gamma$ point. The intra-unit cell interaction between these local AAp orbitals, $U\sim$ 32.6 meV, is much larger than their bandwidth. These orbitals are strongly correlated [28, 19] and tend to form local moments. The $lc$ orbitals are less correlated. The model resembles an extended heavy-fermion like model with a strongly ${\bf k}$-dependent hybridization between the local AAp and the itinerant $lc$ orbitals, concentrated around $\Gamma$, see Fig. S1 [45] and [18, 28, 29, 19, 49, 50, 21, 20, 51]. The interactions involving the itinerant orbitals are of order $U/2$ and cannot be neglected [45, 28, 19, 1, 29]. As in our previous work, we treat the intra and inter-orbital onsite interaction $U$ between the AAp orbitals within DMFT and the other interactions at a mean field Hartree level, within a double self-consistency loop (DMFT+H), see [19, 45]. Temperature is $5.8\,\rm K$ except otherwise indicated. Spontaneous symmetry-breaking is not allowed in the calculation.
@@
Consistent with our previous work [19] and with experimental results [1, 5], the DOS and the inverse compressibility in Fig. 1(a)-(b), respectively, show cascades of spectral weight and asymmetric saw-tooth peaks resetting at integer values of doping. Regions of negative inverse compressibility appear below the integer values after the contribution of the gate is taken into account [45, 20], in agreement with experimental data. The remote bands approach the chemical potential around $\Gamma$, but without crossing the chemical potential at dopings $|\nu|<4$, see Fig. 3(d)-(e) and [45]. Consequently, the system is insulating at $|\nu|=4$, instead of metallic as in [19]. In the present work we have increased $\Delta$, the energy gap between the remote and the flat bands at $\Gamma$, and the dielectric constant $\epsilon$, reducing, therefore, the value of the interactions, as compared with  [19]. While these changes imply a reduction of the ratio between $U$ and $\Delta$ from $U/\Delta=$2 in  [19] to $U/\Delta=$0.85 here, we note that the extent to which the remote bands approach the flat bands, and whether they cross the chemical potential or not, is not simply controlled by the ratio $U/\Delta$ but involves the difference in the interactions which only affect the local AAp electrons and those involving in some way the itinerant ones. The larger value of $\epsilon$ used here reduces also this relative interaction energy.